OpenStudy (anonymous):

The derivative of h(x) is given by the function h '(x) = x + 2cosx. Use your calculator to determine the number of times the graph of h(x) changes concavity on the interval [0,4pi ]. Justify your answer using a complete sentence.

OpenStudy (anonymous):

\[ h''(x) =1 - 2 \sin(x) \]

OpenStudy (anonymous):

how did you get it?

OpenStudy (anonymous):

what's the derivative of x?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

What's the derivative of 2 cos(x)

OpenStudy (anonymous):

it is -2sin

OpenStudy (anonymous):

What is the derivative of h'(x)=x + 2 cos(x)

OpenStudy (anonymous):

o thats it then?

OpenStudy (anonymous):

yes. Now you have to see where 1 - 2 sin(x) is positive and when it is negative.

OpenStudy (anonymous):

it is positive when \( 0\le x \le \frac \pi 6\)

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

do you have a website with more examples?

OpenStudy (anonymous):

It is also positive for \( \frac {5 \pi} 6 \le x \le 2\pi \)

OpenStudy (anonymous):

it is negative when \( \frac \pi 6 \le x \le \frac {5 \pi} 6\)

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

see my site http://saab.org

OpenStudy (anonymous):

in particular, visit http://moltest.missouri.edu/mucgi-bin/calculus.cgi

OpenStudy (anonymous):

So it changes twice then...right?